OpenStudy (anonymous):

find gradient of the curve y=x3+5x+3 at the point(3,15)

OpenStudy (nubeer):

can you find the derivaative?

OpenStudy (anonymous):

FIND DERIVATIVE WITH RESPECT TO X AT X=3 dy/dx=3x^2+5 AT X=3 dy/dx=3(3)^2+5=27+5=32=GRADIENT OF CURVE AT X=3:)